# Assessing subhalo finders in cosmological hydrodynamical simulations

### v1

@@ 3.2.5 Merging
Pairs of structures with similar masses can merge as a consequence of dynamical friction, whereby the lower mass one loses orbital energy and eventually becomes deeply embedded in the core of its host. This process is sometimes insufficient to disrupt the sinking subhalo  [^5] If the sinking process does not sufficiently disturb the subhalo so as to disrupt it, checking its self-boundness is similar to checking the self-boundness of a subset of the most bound particles of its host halo. Hence, it is likely to be found as self-bound., which for a tracking subhalo finder like HBT-HERONS can result in persisting subhaloes that heavily overlap in phase-space. This is not a problem for other methods of halo finding, since density peaks would coalesce into one and configuration and phase-space finders would in principle not find two separate subhaloes.
@@
$\Delta=\dfrac{|\vec{x}_{i}-\vec{x}_{j}|}{\sigma_{i,x}}+\dfrac{|\vec{v}_{i}-\vec{v}_{j}|}{\sigma_{i,v}}$ (6)
Here, $\sigma_{x}$ and $\sigma_{v}$ are, respectively, the position and velocity dispersion of the 10 most bound tracer particles of the parent subhalo, which are also used to compute its position $\vec{x}_{i}$ and velocity $\vec{v}_{i}$. The position and velocity of the subhalo located deeper in the hierarchy [^6] In HBT-HERONS, depth refers to the number of hierarchical connections that a given subhalo is away from the central subhalo of its host FoF group. A central subhalo thus has a depth of 0, its satellites a depth of 1, the satellites of satellites a depth of 2, etc. are also estimated using its 10 most bound tracer particles.
@@
This condition is checked between pairs of subhaloes contained in the same hierarchical tree branch, e.g. TrackIDs 0, 3 and 1 in output 4 of Fig. 1. If $\Delta\leq 2$, the less massive subhalo is marked as merged, and all of the particles in its source subhalo are acquired by the more massive one. This is followed by a new unbinding of the subhalo that accreted the mass, to update its bound mass based on the gravitational potential that includes the newly accreted mass.
@@
As merger checks can only happen between subhaloes that share an explicit hierarchical connection, not all potential satellite-satellite merger configurations are handled by HBT-HERONS. For example, the current approach is unable to account for mergers between satellite subhaloes that were independently accreted as separate centrals, but subsequently encounter each other at sufficiently low velocities so as to form a bound system. The fact that the subhaloes were accreted separately means that no hierarchical relation links them, and hence no merger checks occur between them. Future work will modify HBT-HERONS to be able to handle such encounters.
@@ 3.4 ROCKSTAR
Contrary to the other subhalo finders used in this work, ROCKSTAR only does a single unbinding iteration, and keeps all subhaloes whose 6D FoF group contains more than a specified fraction of bound particles (0.5 by default). This approach is computationally cheap, although an iterative unbinding approach is able to cut down the number of spurious subhaloes and identify otherwise unresolved ones, as discussed in Griffen et al. (2016). In this work, we use the default of a single unbinding pass, to reflect the choice made in other large-scale cosmological simulations, such as the Euclid flagship simulations (Euclid Collaboration et al., 2024b). Another difference relative to the other three subhalo finders is that ROCKSTAR includes particles that belong to parent subhaloes in the calculation of particle gravitational potentials if the subhalo being analysed is undergoing a major merger. The most important difference for the purposes of this study is that ROCKSTAR allows particles to be bound to multiple subhaloes at once (i.e. it uses an inclusive mass assignment), whereas the other three subhalo finders only allow any given particle to be bound to one subhalo (i.e. exclusive mass assignment). We outline how we address this difference in §3.6.
@@ 3.6 Halo and subhalo properties
To do so, we modified ROCKSTAR so that it assigns masses to subhaloes exclusively, similar to what is done in the other subhalo finders we compare in this work. We also compute properties for HBT-HERONS, Subfind and VELOCIraptor subhaloes using SOAP [^8] Available at https://github.com/SWIFTSIM/SOAP (Spherical Overdensity and Aperture Processor; McGibbon et al. in prep.). SOAP is a Python package that computes a variety of subhalo properties in different spherical and projected apertures, both with and without taking into account the bound membership of particles. In short, SOAP loads the centres of subhaloes and memberships of particles as given by each subhalo finder [^9] ROCKSTAR does not provide the bound membership of particles in its outputs, meaning that we only use SOAP for the values of $M_{\mathrm{200c}}$. However, bound memberships are only required for bound mass and maximum circular velocity calculations. We have verified that the relevant routines in SOAP and ROCKSTAR are identical., and computes the properties of interest using the same routines, hence circumventing additional choices made by each particular halo finding algorithm. We explore how the changes we made to the mass assignment in ROCKSTAR and how SOAP-computed $M_{200\mathrm{c}}$ values differ from those provided by subhalo finders in Appendix E.
@@ 4.3.1 Number of subhaloes in the outskirts of central subhaloes
We start the comparison by looking at how many subhaloes there are near $R_{\mathrm{200c}}$ of central subhaloes, measured using a bin of width 0.1 dex. This corresponds to approximately $[0.8-1]\times R_{\mathrm{200c}}$. Our focus on the outskirts of haloes reflects the fact that we expect the difficulties associated with heavily overlapping or disturbed subhaloes to be less important than in the inner regions. Thus, this test will reveal whether there are already systematic differences in the number of subhaloes in regions expected to be relatively well resolved.
Figure 9: Top panels: Mean number density of subhaloes in the outskirts of haloes, as a function of their central $M_{\mathrm{200c}}$ mass. The number density is measured by counting subhaloes with a total bound mass greater than 100 DMO dark matter particles in a spherical shell spanning $[0.8-1]\times R_{\mathrm{200c}}$ of the corresponding host. All subhaloes within the shell are included regardless of whether they are a central or a satellite, in order to minimise differences that are driven purely by the definition of a central. Results from the DMO and hydrodynamical versions of the boxes are shown in the left and right columns, respectively. Bottom panels: The same as above, but expressed relative to the number density found by HBT-HERONS. Note that for the purposes of this comparison, we modified ROCKSTAR to assign masses exclusively, which is the way mass is assigned by the other subhalo finders, but it is not ROCKSTAR’s default behaviour.
@@
Figure 10: Mean subhalo number density as a function of normalised distance to the central subhalo for different DMO resolutions (different line styles) and averaged in different bins of $M_{\mathrm{200c}}$ (different line colours and rows). The values of $M_{\mathrm{200c}}$ that bracket each bin are indicated on the right hand size of each corresponding row, and encompass 0.2 dex in $M_{\mathrm{200c}}$ and are spaced 0.6 dex from each other. Each column corresponds to the results found by a different subhalo finder. The number densities are normalised to the value found by each subhalo finder in the outermost radial shell for each resolution and mass bin. We count all subhaloes whose bound mass is at least 100 times the DMO particle mass, regardless of whether they are classified as centrals or satellites. This choice is made to minimise differences that are driven purely by the definition of a central. Note that for the purposes of this comparison, we modified ROCKSTAR to assign masses exclusively, which is the way mass is assigned by the other subhalo finders, but it is not ROCKSTAR’s default behaviour.
@@ 4.3.2 Radial distributions as a function of host halo mass
Another important property of subhalo populations is how they are distributed within their host haloes. In this subsection, we measure their distribution by counting how many subhaloes are present within a given spherical shell. We use 20 shells whose edges are logarithmically spaced between $[0.01-1]\times R_{\mathrm{200c}}$. A subhalo is assigned to a shell based on where its centre is located. We then divide by the volume of the shell to obtain the subhalo number density as a function of distance to the centre of the host.
As the satellite distribution of an individual halo can be noisy and not representative of the mean trend, we average the number density distributions in nine $M_{\mathrm{200c}}$ bins between $10^{13}$ and $10^{15}\,\,\mathrm{M}_{\odot}$. All but the highest $M_{\mathrm{200c}}$ bin span 0.2 decades in mass. The highest mass bin is enlarged to to 0.4 decades to increase the number of centrals it contains to about 755111111There are small differences in the number of centrals across subhalo finders in the highest $M_{\mathrm{200c}}$ bin due to different $M_{\mathrm{200c}}$ being measured for the same haloes (e.g. Fig 3)..
In the preceding subsection, we found that the number density of subhaloes differs between subhalo finders even near $R_{\mathrm{200c}}$. This means that differences in the distributions will appear due to differences in the normalisation. In this subsection, we are more interested in how the shape of the distribution changes between finders. Therefore, to decouple the effect of different numbers of subhaloes found across subhalo finders, we normalise the subhalo number densities by the number density found by each subhalo finder in the outermost spherical shell (spanning $[0.8-1]\times R_{\mathrm{200c}}$).
Figure 11: Similar to Fig. 10, but shown for the L1_m9 DMO and hydrodynamical simulations and for a larger range of $M_{200\mathrm{c}}$ halo masses. Top panels: the radial distributions found in the DMO version of the simulations. The secondary panel shows the ratio of the normalised subhalo number densities of each subhalo finder relative to the values found in HBT-HERONS Bottom panels: Same as the top panels, but for the hydrodynamical version of the simulations.
@@
Figure 12: Similar to Fig. 11, but the subhalo number density distributions are shown for bins of different satellite-to-host bound mass ratios (different line colours) in the L1_m9 simulation measured for haloes with $M_{200\mathrm{c}}\in[10^{14.6},10^{15.0}]\,\,\mathrm{M}_{\odot}$. Top panels: the radial distributions found in the DMO version of the simulations. The secondary panel shows the ratio of the normalised subhalo number densities of each subhalo finder relative to the values found in HBT-HERONS. Bottom panels: Same as the top panels, but for the hydrodynamical version of the simulations.
@@
Figure 13: Similar to Fig. 12, but the subhalo number density distributions are now shown for across the three available DMO resolutions (different line styles). Different bins of satellite-to-host bound mass ratios are shown using different colours, as indicated in the colour bar. Note that the L1_m10 simulation has a very undersampled distribution in the lowest bin in $M_{\mathrm{satellite}}/M_{\mathrm{central}}$. For our cut of 100 bound particles, $M_{\mathrm{satellite}}\approx 10^{12}\,\mathrm{M}_{\odot}$. This means only very few satellites will contribute to our measurements, as the bin itself requires $M_{\mathrm{satellite}}\leq 10^{-3}\times[10^{14.6},10^{15}]\,\,\mathrm{M}_{\odot}=[10^{11.6},10^{12}]\,\,\mathrm{M}_{\odot}$.
@@ 5 Conclusions
As each finder makes different assumptions and uses different methods to compute subhalo properties, we used an external tool (SOAP; McGibbon et al. in prep) to measure subhalo properties in a consistent manner. Additionally, we also implemented exclusive mass assignment in ROCKSTAR (i.e. a particle can only be bound to a single subhalo) so that it more closely resembles the manner in which the other three subhalo finders assign masses to subhaloes. To avoid trivial resolution effects, we only consider subhaloes with masses greater than 100 dark matter particles. Our main results can be summarised as follows:
@@
    Following from iv), an implicit choice made when running subhalo finders that assign a single central per FoF group (HBT-HERONS, Subfind and VELOCIraptor) is the FoF linking length. Given that there is no a priori correct choice when it comes to comparisons to observations, we explored how strongly the choice of linking length affects the resulting z = 0 $M_{\mathrm{200c}}$ mass functions of HBT-HERONS. We did this by varying the fiducial linking length of 0.2 to 0.16 and 0.25 times the mean interparticle separation. These linking lengths roughly correspond to a factor of two lower (0.25) and higher (0.16) overdensities relative to the commonly used choice. We found steeper and shallower mass functions for the shorter and longer linking lengths, respectively. This can change the predicted abundance of Milky Way-mass haloes by up to 6%, and the difference only drops below 1% for the most massive clusters of galaxies (Fig. 4).
@@
    The bound mass functions differ by larger amounts ($\approx 50\%$) than the $M_{\mathrm{200c}}$ mass functions (Fig. 6). Subfind assigns less mass to satellites than HBT-HERONS, which boosts the bound mass of the central subhaloes it finds. VELOCIraptor and HBT-HERONS agree well, although there are slightly more massive satellites in VELOCIraptor. This is likely due to the central definition, as VELOCIraptor does not require the most massive subhalo to be the central of a FoF group. ROCKSTAR satellites have a lot less bound mass than for the other finders, and the associated number densities also change more strongly with resolution (Fig. 7). Differences remain even when counting subhaloes as a function of $V_{\mathrm{max}}$ instead of $M_{\mathrm{bound}}$ (Fig. 8), although it improves the agreement between Subfind and HBT-HERONS.
@@
    The normalisation and shape of the subhalo radial distributions are both sensitive to which subhalo finder is used. Close to $R_{\mathrm{200c}}$, where the difficulties associated with analysing the dense core of haloes are less important, the number density of subhaloes, $n(r)$, can vary by as much as $20\%$ between subhalo finders (Fig. 9). The shapes of the distribution within $R_{\mathrm{200c}}$ can also change dramatically, with $10\%$ agreement for $n(r)/n(R_{\mathrm{200c}})$ only reached beyond $0.1R_{\mathrm{200c}}$ (Fig. 11). Out of the fours subhalo finders we compare, Subfind struggles the most to find subhaloes near the central regions of haloes, as it exhibits a clear suppression in the number density of subhaloes at small radii regardless of $M_{200\mathrm{c}}$. The problem with the identification of subhaloes by Subfind worsens as the mass of the satellite subhalo relative to its host increases, meaning that either the most massive satellite subhaloes are typically missing or their masses are severely suppressed relative to the values found in the other subhalo finders (Fig. 12). This occurs because the chosen central subhalo acquires most of the mass in the FoF group, leaving a subhalo with an otherwise comparable mass, severally depleted in particle content.
@@
    The inclusion of hydrodynamics increases the number of subhaloes near $R_{\mathrm{200c}}$ and steepens their number density profiles. This is likely driven by the cores of subhaloes becoming denser, both due to baryons being more concentrated than the DM and the increase in the central DM density in response to the baryons. As denser subhaloes are more resilient to disruption, more survive to $z=0$. Additionally, the agreement between subhalo finders tends to improve, particularly for Subfind. This is because the more prominent density peaks are easier to identify against the backdrop of the halo they are embedded in.
@@
    Satellites are segregated by their mass relative to that of the host, with the more massive ones being more concentrated towards the centre of the host halo (Fig. 12). This is opposite to the trend found in previous studies based on Subfind, which claimed that more massive satellites are less concentrated towards the centre of their host. We think the interpretation offered in those studies suffered from systematically biased subhalo catalogues. This major discrepancy across subhalo finders remains across resolution levels (Fig. 13). In other words, different subhalo finders converge to different answers.
@@
In short, (sub)halo finding is not a solved problem, and the choice of which algorithm to use needs to be made carefully. As it is unlikely, and probably undesirable, for the whole community to adopt a single approach to finding subhaloes and to use the same definitions – and none of the (sub)halo finding algorithms can be directly applied to observations – part of the uncertainty budget in making predictions should be devoted to the manner in which (sub)haloes are found. A related, but slightly different aspect, is how the properties of galaxies change across subhalo finders. As they are located in the centres of dark matter subhaloes, and the identification of a subhalo core is typically easier than the identification of its edge, the properties of galaxies may not differ as much as what we have discussed in this study.

### v3

@@ 3.2.5 Merging
Pairs of structures with similar masses can merge as a consequence of dynamical friction, whereby the least massive subhalo loses orbital energy and eventually becomes deeply embedded in the core of its host. This process is sometimes insufficient to fully disrupt the sinking subhalo  [^7] If the sinking process does not sufficiently disturb the subhalo so as to disrupt it, checking its self-boundness is similar to checking the self-boundness of a subset of the most bound particles of its host halo. Hence, it is likely to be found as self-bound even though it no longer represents a separate physical structure., which for a tracking subhalo finder like HBT-HERONS can result in persisting subhaloes that heavily overlap in phase-space. This is not a problem for other methods of subhalo finding, since density peaks would coalesce into one and configuration and phase-space finders would in principle not find two separate subhaloes.
@@
$\Delta_{ij}=\dfrac{|\vec{x}_{i}-\vec{x}_{j}|}{\sigma_{i,x}}+\dfrac{|\vec{v}_{i}-\vec{v}_{j}|}{\sigma_{i,v}}$ (6)
Here, $\sigma_{i,x}$ and $\sigma_{i,v}$ are, respectively, the position and velocity dispersion of the $N^{\rm min}_{\rm core}$ (20 by default) most bound tracer particles of subhalo $i$, which are also used to compute its position $\vec{x}_{i}$ and velocity $\vec{v}_{i}$. The position and velocity of subhalo $j$ are also estimated using its $N^{\rm min}_{\rm core}$ most bound tracer particles. The number of tracer particles used to estimate the properties of the subhalo core should be $N^{\rm min}_{\rm core}\approx N^{\rm min}_{\rm bound}$. Using a value much larger than $N^{\rm min}_{\rm bound}$ means that the measured phase-space dispersion is no longer representative of the core of the subhalo, but of the whole subhalo. This can lead to premature merging between pairs of well resolved subhaloes that are neither nearby in space nor stationary with respect to each other.
We compute the phase-space offset in both directions, i.e. one time using the measured dispersion of subhalo $i$ and another time using the dispersion of subhalo $j$, and take the minimum of the two values: $\Delta_{\rm min}=\mathrm{min}(\Delta_{ij},\Delta_{ji})$. Usually, the more massive subhalo of the pair has a larger phase-space dispersion than the less massive one, but this is not necessarily the case when both subhaloes are relatively poorly resolved ($N_{\mathrm{bound}}\lesssim 100$).
@@
This condition is checked between pairs of subhaloes contained in the same hierarchical tree branch, e.g. TrackIDs 0, 3 and 1 in output 4 of Fig. 1. If $\Delta_{\rm min}\leq 2$, the subhalo located deeper in the hierarchy is marked as merged, and all of the particles in its source subhalo are acquired by the more massive one. This is followed by a new unbinding of the subhalo that accreted the mass, to update its bound mass based on the gravitational potential that includes the newly accreted mass. Consequently, both central and satellite subhaloes are eligible to grow in mass if they overlap with a subhalo contained in their own satellite system.
@@
As merger checks can only happen between subhaloes that share an explicit hierarchical connection, not all potential satellite-satellite merger configurations are handled by HBT-HERONS. For example, the current approach is unable to account for mergers between satellite subhaloes that were independently accreted as separate centrals, but subsequently encounter each other at sufficiently low velocities so as to form a bound system. The fact that the subhaloes were accreted separately means that no hierarchical relation links them, and hence no merger checks occur between them. Nonetheless, this is minute fraction of satellite-satellite mergers, which are themselves not very common (Bahé et al., 2019). Future work will modify HBT-HERONS to be able to handle such encounters.
3.2.6 Non-hierarchical transfer of mass
An important caveat to the mass transfer channels discussed above is that they are implicitly designed to be hierarchical in nature, e.g. low mass subhaloes can only transfer mass to more massive subhaloes, and not vice versa. This underlying assumption works well for particles that are collisionless in nature, like dark matter or stars. However, gas is inherently collisional, meaning that the assumption of purely hierarchical mass transfer may not work as well in simulations that contain gas particles. Not accounting for the possibility of non-hierarchical gas exchange would entail that gas particles (or eventually star particles formed from them) will (incorrectly) remain assigned to their original subhalo, rather than the one in which they now physically reside.
To avoid this problem, we perform a re-attachment of gas particles that does not assume anything about the underlying hierarchical connections between subhaloes, and hence allows for the transfer of mass between arbitrary subhaloes within a FoF groups (as long as the local physical conditions are conducive for it to happen). In short, for each gas particle in a FoF group we find their 10 nearest tracer particles. We then check whether any of those neighbours belong to the same TrackID as the gas particle itself. If none do, we remove the gas particle from the source subhalo of its original TrackID, and re-attach it to the source subhalo of the TrackID with the largest representation among the 10 nearest tracer particles. This approach means that only gas particles that are truly far away from their supposed subhalo are allowed to be re-attached. In practice, this step is done before before any subhalo is subject to unbinding, so that the re-attached gas can become bound to their newly assigned TrackID. More details about the re-attachment are provided in §A.8
@@ 3.4 ROCKSTAR
Contrary to the other subhalo finders used in this work, ROCKSTAR only does a single unbinding iteration, and keeps all subhaloes whose 6D FoF group contains more than a specified fraction of bound particles (0.5 by default). This approach is computationally cheap, although an iterative unbinding approach is able to cut down the number of spurious subhaloes and identify otherwise unresolved ones, as discussed in Griffen et al. (2016). In this work, we use the default of a single unbinding pass, to reflect the choice made in other large-scale cosmological simulations, such as the Euclid flagship simulations (Euclid Collaboration et al., 2025b). Another difference relative to the other three subhalo finders is that ROCKSTAR includes particles that belong to parent subhaloes in the calculation of particle gravitational potentials if the subhalo being analysed is undergoing a major merger. The most important difference for the purposes of this study is that ROCKSTAR allows particles to be bound to multiple subhaloes at once (i.e. it uses an inclusive mass assignment), whereas the other three subhalo finders only allow any given particle to be bound to one subhalo (i.e. exclusive mass assignment). We outline how we address this difference in §3.6.
@@ 3.6 Halo and subhalo properties
To do so, we modified ROCKSTAR so that it assigns masses to subhaloes exclusively, similar to what is done in the other subhalo finders we compare in this work. To emphasise the difference in mass assignment of this modified version of ROCKSTAR from its default, we refer to it as ROCKSTAR (Exc.) whenever it is relevant to the plots being shown. We also compute properties for HBT-HERONS, Subfind and VELOCIraptor subhaloes using SOAP [^9] Available at https://github.com/SWIFTSIM/SOAP (Spherical Overdensity and Aperture Processor; McGibbon et al. 2025). SOAP is a Python package that computes a variety of subhalo properties in different spherical and projected apertures, both with and without taking into account the bound membership of particles. In short, SOAP loads the centres of subhaloes and memberships of particles as given by each subhalo finder [^10] ROCKSTAR does not provide the bound membership of particles in its outputs, meaning that we only use SOAP for the values of $M_{\mathrm{200c}}$. However, bound memberships are only required for bound mass and maximum circular velocity calculations. We have verified that the relevant routines in SOAP and ROCKSTAR are identical., and computes the properties of interest using the same routines, hence circumventing additional choices made by each particular halo finding algorithm. We explore how the changes we made to the mass assignment in ROCKSTAR and how SOAP-computed $M_{200\mathrm{c}}$ values differ from those provided by subhalo finders in Appendix E.
@@ 4.3.1 Number of subhaloes in the outskirts of central subhaloes
We start the comparison by looking at how many subhaloes there are near $R_{\mathrm{200c}}$ of central subhaloes, measured using a bin of width 0.1 dex. This corresponds to approximately $0.8-1]\times R_{\mathrm{200c}}$. Our focus on the outskirts of haloes reflects the fact that we expect the difficulties associated with heavily overlapping or disturbed subhaloes to be less important than in the inner regions. Thus, this test will reveal whether there are already systematic differences in the number of subhaloes in regions expected to be relatively well resolved.
Figure 9: Top panels: Mean number density of subhaloes in the outskirts of haloes, as a function of their central $M_{\mathrm{200c}}$ mass. The number density is measured by counting subhaloes with a total bound mass greater than 100 DMO dark matter particles in a spherical shell spanning $0.8-1]\times R_{\mathrm{200c}}$ of the corresponding host. All subhaloes within the shell are included regardless of whether they are a central or a satellite, in order to minimise differences that are driven purely by the definition of a central. Results from the DMO and hydrodynamical versions of the boxes are shown in the left and right columns, respectively. Bottom panels: The same as above, but expressed relative to the number density found by HBT-HERONS. Note that for the purposes of this comparison, we modified ROCKSTAR to assign masses exclusively, which is the way mass is assigned by the other subhalo finders, but it is not ROCKSTAR’s default behaviour.
@@
Figure 10: Normalised subhalo number density as a function of normalised distance to the central subhalo for different DMO resolutions (different line styles) and averaged in different bins of $M_{\mathrm{200c}}$ (different line colours and rows). The values of $M_{\mathrm{200c}}$ that bracket each bin are indicated on the right hand size of each corresponding row, and encompass 0.2 dex in $M_{\mathrm{200c}}$ and are spaced 0.6 dex from each other. Each column corresponds to the results found by a different subhalo finder. The number densities are normalised to the value found by each subhalo finder in the outermost radial shell for each resolution and mass bin. We count all subhaloes whose bound mass is at least 100 times the DMO particle mass, regardless of whether they are classified as centrals or satellites. This choice is made to minimise differences that are driven purely by the definition of a central. Note that for the purposes of this comparison, we modified ROCKSTAR to assign masses exclusively, which is the way mass is assigned by the other subhalo finders, but it is not ROCKSTAR’s default behaviour.
@@ 4.3.2 Radial distributions as a function of host halo mass
Another important property of subhalo populations is how they are distributed within their host haloes. In this subsection, we measure their distribution by counting how many subhaloes are present within a given spherical shell. We use 20 shells whose edges are logarithmically spaced between $0.01-1]\times R_{\mathrm{200c}}$. A subhalo is assigned to a shell based on where its centre is located. We then divide by the volume of the shell to obtain the subhalo number density as a function of distance to the centre of the host.
As the satellite distribution of an individual halo can be noisy and not representative of the mean trend, we average the number density distributions in nine $M_{\mathrm{200c}}$ bins between $10^{13}$ and $10^{15}\,\,\mathrm{M}_{\odot}$. All but the highest $M_{\mathrm{200c}}$ bin span 0.2 decades in mass. The highest mass bin is enlarged to to 0.4 decades to increase the number of centrals it contains to about 755121212There are small differences in the number of centrals across subhalo finders in the highest $M_{\mathrm{200c}}$ bin due to different $M_{\mathrm{200c}}$ being measured for the same haloes (e.g. Fig 3)..
In the preceding subsection, we found that the number density of subhaloes differs between subhalo finders even near $R_{\mathrm{200c}}$. This means that differences in the distributions will appear due to differences in the normalisation. In this subsection, we are more interested in how the shape of the distribution changes between finders. Therefore, to decouple the effect of different numbers of subhaloes found across subhalo finders, we normalise the subhalo number densities by the number density found by each subhalo finder in the outermost spherical shell (spanning $0.8-1]\times R_{\mathrm{200c}}$).
Figure 11: Normalised subhalo number density as a function of normalised distance to the central subhalo for the L1_m9 DMO and hydrodynamical simulations averaged in different bins of $M_{\mathrm{200c}}$ (different line colours). Top panels: the radial distributions found in the DMO version of the simulations. The secondary panel shows the ratio of the normalised subhalo number densities of each subhalo finder relative to the values found in HBT-HERONS Bottom panels: Same as the top panels, but for the hydrodynamical version of the simulations.
@@
Figure 12: Normalised subhalo number density as a function of normalised distance to central subhalos with $M_{200\mathrm{c}}\in[10^{14.6},10^{15.0}]$, averaged in different satellite-to-host bound mass ratios (different line colours) for the L1_m9 DMO and hydrodynamical simulations Top panels: the radial distributions found in the DMO version of the simulations. The secondary panel shows the ratio of the normalised subhalo number densities of each subhalo finder relative to the values found in HBT-HERONS. Bottom panels: Same as the top panels, but for the hydrodynamical version of the simulations.
@@
Figure 13: Similar to Fig. 12, but the subhalo number density distributions are now shown across the three available DMO resolutions (different line styles). Different bins of satellite-to-host bound mass ratios are shown using different colours, as indicated in the colour bar. Note that the L1_m10 simulation has a very undersampled distribution in the lowest bin in $M_{\mathrm{satellite}}/M_{\mathrm{central}}$. For our cut of 100 bound particles, $M_{\mathrm{satellite}}\approx 10^{12}\,\mathrm{M}_{\odot}$. This means only very few satellites will contribute to our measurements, as the bin itself requires $M_{\mathrm{satellite}}\leq 10^{-3}\times[10^{14.6},10^{15}]\,\,\mathrm{M}_{\odot}=[10^{11.6},10^{12}]\,\,\mathrm{M}_{\odot}$.
@@ 5 Conclusions
As each finder makes different assumptions and uses different methods to compute subhalo properties, we used an external tool (SOAP; McGibbon et al. 2025) to measure subhalo properties in a consistent manner. Additionally, we also implemented exclusive mass assignment in ROCKSTAR (i.e. a particle can only be bound to a single subhalo) so that it more closely resembles the manner in which the other three subhalo finders assign masses to subhaloes. To avoid trivial resolution effects, we only consider subhaloes with masses greater than 100 dark matter particles. Our main results can be summarised as follows:
@@
    Following from iv), an implicit choice made when running subhalo finders that assign a single central per FoF group (HBT-HERONS, Subfind and VELOCIraptor) is the FoF linking length. Given that there is no a priori physically correct choice when it comes to comparisons to observations, as it is impossible to run a 3D FoF algorithm on the underlying DM distribution around observed galaxies, we explored how strongly the choice of linking length affects the resulting z = 0 $M_{\mathrm{200c}}$ mass functions of HBT-HERONS. We did this by varying the fiducial linking length of 0.2 to 0.16 and 0.25 times the mean interparticle separation. These linking lengths roughly correspond to a factor of two lower (0.25) and higher (0.16) overdensities relative to the commonly used choice. We found steeper and shallower mass functions for the shorter and longer linking lengths, respectively. This can change the predicted abundance of Milky Way-mass haloes by up to 6%, and the difference only drops below 1% for the most massive clusters of galaxies (Fig. 4).
@@
    The bound mass functions differ by larger amounts ($\approx 50\%$) than the $M_{\mathrm{200c}}$ mass functions (Fig. 6). Subfind assigns less mass to satellites than HBT-HERONS, a well-known shortcoming which boosts the bound mass of the central subhaloes it finds. VELOCIraptor and HBT-HERONS agree well, although there are slightly more massive satellites in VELOCIraptor. This is likely due to the central definition, as VELOCIraptor does not require the most massive subhalo to be the central of a FoF group. ROCKSTAR satellites have a lot less bound mass than for the other finders, and the associated number densities also change more strongly with resolution (Fig. 7). Differences remain even when counting subhaloes as a function of $V_{\mathrm{max}}$ instead of $M_{\mathrm{bound}}$ (Fig. 8), although it improves the agreement between Subfind and HBT-HERONS.
@@
    The normalisation and shape of the subhalo radial distributions are both sensitive to which subhalo finder is used. Close to $R_{\mathrm{200c}}$, where the difficulties associated with analysing the dense core of haloes are less important, the number density of subhaloes, $n(r)$, can vary by as much as $20\%$ between subhalo finders (Fig. 9). The shapes of the distribution within $R_{\mathrm{200c}}$ can also change dramatically, with $10\%$ agreement for $n(r)/n(R_{\mathrm{200c}})$ only reached beyond $0.1R_{\mathrm{200c}}$ (Fig. 11). Out of the fours subhalo finders we compare, Subfind struggles the most to find subhaloes near the central regions of haloes, as it exhibits a clear suppression in the number density of subhaloes at small radii regardless of $M_{200\mathrm{c}}$. The problem with the identification of subhaloes by Subfind worsens as the mass of the satellite subhalo relative to its host increases, meaning that either the most massive satellite subhaloes are typically missing or their masses are severely suppressed relative to the values found in the other subhalo finders (Fig. 12). This occurs because the a significant fraction of the satellite mass is artificially assigned to the chosen central subhalo, leaving a subhalo with an otherwise comparable mass severally depleted in particle content.
@@
    The inclusion of hydrodynamics increases the number of subhaloes near $R_{\mathrm{200c}}$ and steepens their number density profiles. This is likely driven by the cores of subhaloes becoming denser, both due to baryons being more concentrated than the DM and the increase in the central DM density in response to the baryons. As denser subhaloes are more resilient to disruption and stripping, more survive to $z=0$. Additionally, the agreement between subhalo finders tends to improve, particularly for Subfind. This is because the more prominent density peaks are easier to identify against the backdrop of the halo they are embedded in.
@@
    Satellites are segregated by their mass relative to that of the host, with the more massive ones being more concentrated towards the centre of the host halo (Fig. 12). This is opposite to the trend found in previous studies based on Subfind, which claimed that more massive satellites are less concentrated towards the centre of their host. We think the interpretation offered in those studies suffered from systematically biased subhalo catalogues. This major discrepancy across subhalo finders remains across resolution levels (Fig. 13). In other words, different subhalo finders converge to qualitatively different answers.
@@
In short, (sub)halo finding remains an open problem, and the choice of which algorithm to use needs to be made carefully. As it is unlikely, and probably undesirable, for the whole community to adopt a single approach to finding subhaloes and to use the same definitions – and none of the (sub)halo finding algorithms can be directly applied to observations – part of the uncertainty budget in making predictions should be devoted to the manner in which (sub)haloes are found. A related, but slightly different aspect, is how the properties of galaxies change across subhalo finders. As they are located in the centres of dark matter subhaloes, and the identification of a subhalo core is typically easier than the identification of its edge, the properties of galaxies may not differ as much as what we have discussed in this study.